# Rethinking Patch Dependence for Masked Autoencoders

### v1

@@ 3.2 Reconstruction with Cross-Attention
To address the first question, our method substitutes the self-attention mechanism in the decoder blocks with cross-attention. Specifically, the decoder employs multi-head cross-attention where the queries are the output from previous decoder blocks (or the sum of position embedding of the masked patches and mask token for the first decoder block.) The keys and values are derived from the encoded features.
@@
Diverging from the original transformer architecture [56], our decoder omits the precursory causal self-attention stage before the introduction of multi-head cross-attention. This elimination, coupled with the fact that layer normalization and residual connections are only applied along the feature axis but not the token axis, enables the independent decoding of tokens. This design choice is evaluated in the ablation study section to determine its impact on performance.
Given the disparity in the dimensions of the encoder and decoder, MAE adapts the visible features to the decoder’s latent space using a Multilayer Perceptron (MLP). However, in CrossMAE, as encoder features are integrated at various decoder blocks, we embed the projection within the multi-head cross-attention module.
This approach also mirrors aspects of Perceiver IO [33], where cross-attention is employed to amalgamate features pertinent to the task at hand. However, in contrast, we do not restrict the architecture to a single cross-attention unit. In our method, the masked tokens are decoded in a manner more akin to the traditional transformer paradigm [56].
@@ 3.3 Partial Reconstruction
The fact that CrossMAE uses cross-attention rather than self-attention in the decoder blocks brings an additional benefit over the original MAE architecture. Recall that mask tokens are decoded independently and thus there is no exchange of information between them, to obtain the reconstructions at a specific spatial location, CrossMAE only needs to pass the corresponding mask tokens to the cross-attention decoder. This allows efficient partial reconstruction in contrast to the original MAE architecture which needs to pass all the masked tokens as the input of the decoder blocks due to the existence of self-attention in the decoder blocks.
To address the second question in Sec. 3.1, rather than decoding the reconstruction for all masked locations, we only compute the reconstruction on a random subset of the locations and apply the loss to the decoded locations. Specifically, we name the ratio of predicted tokens to all image tokens as prediction ratio ($\gamma$), and the mask ratio ($p$). Then the prediction ratio is bounded between $\gamma\in(0,p]$. Because we are sampling within the masked tokens uniformly at random and the reconstruction loss is a mean square error on the reconstructed patches, the expected loss is the same as in MAE, while the variance is ($p/\gamma$) times larger than the variance in MAE. Empirically, we find that scaling the learning rate of MAE ($\beta$) to match the variance (i.e. setting the learning rate as $\gamma\beta/p)$) helps with model performance. Since cross-attention has linear complexity with respect to the number of masked tokens, this partial reconstruction paradigm leads to a decrease in computation complexity. Empirically, we find that the quality of the learned representations is not compromised by this approach.
@@ 3.4 Inter-block Attention
MAE combines the feature of the last encoder block with the mask tokens as the input to the self-attention decoder. This design poses the latent feature as an information bottleneck. Concretely, it leaves no room for any residual connection for the decoder to leverage early encoder features since the decoder blocks sequentially update and improve on the decoded features. In contrast, the cross-attention decoder in CrossMAE decouples queries from keys and values. This adds flexibility as different cross-attention blocks can take different feature maps as keys and values and thus no longer need to only use features from the last encoder block.
Besides simply using the feature from the last encoder block, one naive choice is to give the feature of the $i$th encoder block to the last $i$th decoder (e.g., feed the feature of the first encoder to the last decoder), in a U-Net-like fashion. However, this assumes the depth of the decoder to be the same as the depth of the encoder, which does not hold in MAE where an asymmetric design is adopted.
Instead of manually selecting the feature for each decoder block, we follow our design choice of using cross-attention blocks for inter-patch spatial cross-attention and propose inter-block attention for feature fusion for each decoder block 4(c). Analogous to the inter-patch cross-attention that takes a weighted sum of the visible token embeddings across the patch dimensions to update the embeddings of masked tokens, inter-block attention takes a weighted sum of the visible token embeddings across different input blocks at the same spatial location to fuse the input features from multiple blocks into one feature map for each decoder block.
@@
In addition to feature maps from different encoder blocks, we also include the inputs to the first encoder block to allow the decoder to leverage more low-level information to reconstruct the original image. We can select a subset of the feature maps from the encoder layers instead of all feature maps. This reduces the computation complexity of the system. We ablate over this design choice in LABEL:tab:ablation-wfm.
We show that using the weighted features rather than simply using the features from the last block greatly improves the performance of our CrossMAE. Interestingly, as we will show in Sec. 4.4, early decoder blocks focus on the information from the later encoder blocks, and later decoder blocks focus on the information from the early encoder blocks in the process of learning to achieve better reconstructions.
Method Pre-train Data ViT-S ViT-B ViT-L
Supervised [31, 50] - 79.0 82.3 82.6
DINO [10] IN1K - 82.8 -
MoCo v3 [15] IN1K 81.4 83.2 84.1
BEiT [4] IN1K+DALLE - 83.2 85.2
MultiMAE [3] IN1K - 83.3 -
MixedAE [11] IN1K - 83.5 -
CIM [22] IN1K 81.6 83.3 -
MAE [31] IN1K 78.9 83.3 85.4
CrossMAE (25%) IN1K 79.2 83.5 85.4
CrossMAE (75%) IN1K 79.3 83.7 85.4
@@ 4 Experiments
We perform self-supervised pretraining on ImageNet-1K, following MAE [31]. We perform minimal hyperparameter tuning, maintaining consistency with MAE’s parameters except for the learning rate and decoder depth. The hyperparameters were initially determined on ViT-Base and then directly applied to both ViT-Small and ViT-Large. Both CrossMAE and MAE are trained for 800 epochs. Please refer to the supplementary material for implementation details.
@@ 4.1 ImageNet Classification
APbox APmask
Method Pre-train Data ViT-B ViT-L ViT-B ViT-L
Supervised [39] IN1K w/ labels 47.6 49.6 42.4 43.8
MoCo v3 [15] IN1K 47.9 49.3 42.7 44.0
BEiT [5] IN1K+DALLE 49.8 53.3 44.4 47.1
MixedAE [11] IN1K 50.3 - 43.5 -
MAE [39] IN1K 51.2 54.6 45.5 48.6
CrossMAE IN1K 52.1 54.9 46.3 48.8
Results. As shown in Tab. 1, CrossMAE outperforms vanilla MAE trained on the same ViT-B encoder in terms of fine-tuning accuracy. This shows that replacing the self-attention with cross-attention does not degrade the quality of the pre-trained model in terms of downstream classification. CrossMAE outperforms other strong baselines such as DINO [10], MoCo v3 [15], BEiT [4], and MultiMAE [3].
@@ 4.2 Object Detection and Instance Segmentation
Method
[…]
MAE
83.0
CrossMAE
83.3
CrossMAE + Self-Attn
83.3
Mask Ratio
[…]
$65\%$
83.5
$75\%$
83.3
$85\%$
83.3
Pred. Ratio
[…]
$15\%$
83.1
$25\%$
83.2
$75\%$
83.3
# Feature Maps Fused
[…]
1
82.9
3
83.3
6
83.5
12
83.3
Decoder Depth
[…]
1
83.0
4
83.1
8
83.1
12
83.3
Image Resolution
[…]
224
83.2
448
84.6
Table 3: Ablations on CrossMAE. We report fine-tuning performance on ImageNet-1K classification with 400 epochs (i.e., half of the full experiments) with ViT-B/16. MAE performance is reproduced using the official MAE code. Underline indicates the default setting for CrossMAE. Bold indicates the best hyperparameter among the tested ones. $1$ feature map fused (row 1, Table 2d) indicates using only the feature from the last encoder block. We use $25\%$ prediction ratio for both settings in Table 2f to accelerate training.
Setup. We additionally evaluate models pretrained with CrossMAE for object detection and instance segmentation, which require deeper spatial understanding than ImageNet classification. Specifically, we follow ViTDet [39], a method that leverages a Vision Transformer backbone for object detection and instance segmentation. We report box AP for object detection and mask AP for instance segmentation, following MAE [31]. We compare against 4 baselines: supervised pre-training, MoCo-v3 [15], BEiT [5], and MAE [31].
Results. As listed in Tab. 2, CrossMAE, with the default $75\%$ prediction ratio, performs better compared to these strong baselines, including vanilla MAE. This suggests that similar to MAE, CrossMAE performance on ImageNet positively correlates with instance segmentation. Additionally, CrossMAE’s downstream performance scales similarly to MAE as the model capacity increases from ViT-B to ViT-L.
@@ 4.3 Ablations
Cross-Attention vs Self-Attention. As summarized in LABEL:tab:cross-self-results, CrossMAE that uses a cross-attention-only decoder has better downstream performance than vanilla MAE, which is consistent with the primary observations in Sec. 4.1. We further show that applying both cross-attention and self-attention together does not lead to additional benefits in terms of fine-tuning performance compared to using cross-attention only. This suggests that using only cross-attention is sufficient for learning good representations.
Mask Ratio and Prediction Ratio. We ablate over a range of mask ratio (i.e., the ratio of mask tokens to all tokens) and the prediction ratio (i.e., the ratio of reconstructed tokens to all tokens) in LABEL:tab:ablation-mask-ratio and LABEL:tab:ablation-keep-mask-ratio. We observe that our method is not sensitive to varying masked tokens. Furthermore, although predicting the reconstructions from all the mask tokens leads to the best pre-trained model, the difference between full prediction and partial reconstruction is marginal. Specifically, CrossMAE can be trained to reconstruct as few as 15% of the tokens in the decoder rather than 100% of the tokens as required by the vanilla MAE baseline, yet achieving similar downstream finetuning performance. This result suggests that a good representation can be learned by reconstructing only part of an image.
Inter-block Attention. We also vary the number of encoder feature maps that are fused with our inter-block attention as an ablation. In addition to simply taking the feature from the last encoder block (i.e., using only one feature map) and all encoder blocks (i.e., using all 12 feature maps), we uniformly select feature maps to be fused in terms of their encoder block index. As shown in LABEL:tab:ablation-wfm, using only the last feature map leads to a minor degradation of performance compared to using all feature maps. Furthermore, adding even a subset of feature maps boosts the performance of CrossMAE, with the best performance reached when 6 feature maps are fused. This indicates that CrossMAE does not require all feature maps to obtain its optimal performance, which further justifies the efficiency of CrossMAE.
Decoder Depth.
Figure 5: We compare ViT-B which is pre-trained for 800 epochs with different variants of CrossMAE v.s. MAE. For CrossMAE, we vary the prediction ratio $p$ and number of decoder blocks $n$, and we denote each as ($p$, $n$). While all experiments are run with inter-block attention, CrossMAE has lower decoder FLOPS than MAE [31] and performs on par or better.
As shown in LABEL:tab:ablation-decoder-depth, we show that using a decoder of 12 blocks slightly improves downstream performance compared to shallower decoders. CrossMAE performs on par with the vanilla MAE even with as few as one decoder block, which shows CrossMAE’s capability for efficiently pre-train vision transformers. We further conduct full-scale experiments to compare the impact of decoder depth and prediction ratio, as they can both significantly impact the efficiency of the model. Results are summarized in Fig. 5 and the runtimes is compared in Tab. 4. We find that a model trained with a small prediction ratio can benefit more from a deeper decoder.
Figure 6: We visualize the output of each decoder block. (a-b) Different decoder blocks play different roles in the reconstruction, with most details emerging at later decoder blocks, which confirms the motivation for inter-block attention. (c) Visualizations of inter-block attention shows that different decoder blocks indeed attend to feature from different encoder blocks, with later blocks focusing on earlier encoder features to achieve reconstruction.
Method
Pred.
Ratio
Decoder
Depth
Memory Runtime
MAE 0.75 8 OOM ($>$81920) 103.45
CrossMAE 0.25 12 41630 65.80
Table 4: Pretraining runtime and GPU memory comparison on ViT-B for 10 epochs using 2 NVIDIA A100 80GB GPUs. Memory is measured in MB per GPU, and runtime is in minutes. Please refer to Tab. 5 for a complete comparison. MAE trained without using gradient accumulation at the default batch size does not fit onto 2 GPUs, thus the memory usage is not reported.
@@
Input Resolution. We test CrossMAE on longer token lengths by increasing the image resolution without changing the patch size. As the resolution grows from 224 to 448, the image token length increases from 197 to 785, which poses great challenges to the scalability of existing methods. Thus, we deliberately choose the CrossMAE variant with a 25% prediction ratio for higher efficiency. In LABEL:tab:ablation-input-res, we observe that the classification accuracy positively correlates with the input resolution, suggesting that CrossMAE can scale to long input sequences.

### v2

@@ 3.2 Reconstruction with Cross-Attention
To address this question, we substitute the self-attention mechanism in the decoder blocks with cross-attention, using it as a readout function to decode the latent embedding from the encoder to raw pixel values. Specifically, the decoder employs multi-head cross-attention where the queries are the output from previous decoder blocks (or the sum of position embedding of the masked patches and mask token for the first decoder block). The keys and values are from the encoded features.
@@
Diverging from the original transformer architecture (Vaswani et al., 2017), our decoder omits the causal self-attention layer before the introduction of multi-head cross-attention. This elimination, coupled with the fact that layer normalization and residual connections are only applied along the feature axis but not the token axis, enables the independent decoding of tokens. This design choice is evaluated in the ablation study section to determine its impact on performance.
Given the disparity in dimensions between the encoder and decoder, MAE adapts the encoded features to the decoder’s latent space using an MLP. However, in CrossMAE, as encoder features are integrated at various decoder blocks, we embed the projection within the multi-head cross-attention module.
Cross-attention layers serve as a readout function that decodes the global representation provided in the encoder’s output tokens to the pixel values within each patch to be reconstructed. However, CrossMAE does not restrict the architecture to a single cross-attention block. Instead, we stack multiple cross-attention decoder blocks in a manner more akin to the traditional transformer (Vaswani et al., 2017).
@@ 3.3 Partial Reconstruction
The fact that CrossMAE uses cross-attention rather than self-attention in the decoder blocks brings an additional benefit over the original MAE architecture. Recall that mask tokens are decoded independently and thus there is no exchange of information between them, to obtain the reconstructions at a specific spatial location, CrossMAE only needs to pass the corresponding mask tokens to the cross-attention decoder. This allows partial reconstruction in contrast to the original full-image reconstruction in the MAE architecture which needs to pass all the masked tokens as the input of the decoder blocks due to the existence of self-attention in the decoder blocks.
To address the second question in Sec. 3.1, rather than decoding the reconstruction for all masked locations, we only compute the reconstruction on a random subset of the locations and apply the loss to the decoded locations. Specifically, we name the ratio of predicted tokens to all image tokens as prediction ratio ($\gamma$), and the mask ratio ($p$). Then the prediction ratio is bounded between $\gamma\in(0,p]$. Because we are sampling within the masked tokens uniformly at random and the reconstruction loss is a mean square error on the reconstructed patches, the expected loss is the same as in MAE, while the variance is ($p/\gamma$) times larger than the variance in MAE. Empirically, we find that scaling the learning rate of MAE ($\beta$) to match the variance (i.e. setting the learning rate as $\gamma\beta/p)$) helps with model performance. Since cross-attention has linear complexity with respect to the number of masked tokens, this partial reconstruction paradigm decreases computation complexity. Empirically, we find that the quality of the learned representations is not compromised by this approach.
@@ 3.4 Inter-block Attention
MAE combines the feature of the last encoder block with mask tokens as the input to the self-attention decoder, which creates an information bottleneck by making early encoder features inaccessible for the decoder. In contrast, CrossMAE’s cross-attention decoder decouples queries from keys and values. This decoupling allows different cross-attention decoder blocks to take in feature maps from different encoder blocks. This added degree of flexibility comes with a design choice for selecting encoder features for each decoder block. One naive choice is to give the feature of the $i$th encoder block to the last $i$th decoder (e.g., feeding the feature of the first encoder to the last decoder), in a U-Net-like fashion. However, this assumes the decoder’s depth matches the depth of the encoder, which is not the case for MAE or CrossMAE.
Instead of manually matching each decoder block with an encoder feature map, we make the selection learnable and propose inter-block attention for feature fusion for each decoder block (Figure 4(c)). Analogous to the inter-patch cross-attention that takes a weighted sum of the visible token embeddings across the patch dimensions to update the embeddings of masked tokens, inter-block attention takes a weighted sum of the visible token embeddings across different input blocks at the same spatial location to fuse the input features from multiple blocks into one feature map for each decoder block.
@@
In addition to feature maps from different encoder blocks, we also include the inputs to the first encoder block to allow the decoder to leverage more low-level information to reconstruct the original image. We can select a subset of the feature maps from the encoder layers instead of all feature maps. This reduces the computation complexity of the system. We ablate this in Table 3d.
We show that using the weighted features rather than simply using the features from the last block greatly improves the performance of CrossMAE. Intriguingly, in the process of learning to achieve better reconstructions, early decoder blocks tend to prioritize information from later encoder blocks, while later decoder blocks focus on earlier encoder block information, as demonstrated in Section 4.5.
Method ViT-S ViT-B ViT-L ViT-H
Supervised Steiner et al. (2022) 79.0 82.3 82.6 83.1
DINO Caron et al. (2021) - 82.8 - -
MoCo v3 Chen et al. (2021b) 81.4 83.2 84.1 -
BEiT Bao et al. (2021) - 83.2 85.2 -
MultiMAE Bachmann et al. (2022) - 83.3 - -
MixedAE Chen et al. (2023) - 83.5 - -
CIM Fang et al. (2023) 81.6 83.3 - -
MAE He et al. (2022) 78.9 83.3 85.4 85.8
CrossMAE (25%) 79.2 83.5 85.4 86.3
CrossMAE (75%) 79.3 83.7 85.4 86.4
@@
APbox APmask
Method ViT-B ViT-L ViT-B ViT-L
Supervised Li et al. (2022b) 47.6 49.6 42.4 43.8
MoCo v3 Chen et al. (2021b) 47.9 49.3 42.7 44.0
BEiT Bao et al. (2022) 49.8 53.3 44.4 47.1
MixedAE Chen et al. (2023) 50.3 - 43.5 -
MAE Li et al. (2022b) 51.2 54.6 45.5 48.6
CrossMAE 52.1 54.9 46.3 48.8
@@ 4 Experiments
We perform self-supervised pretraining on ImageNet-1K, following MAE (He et al., 2022)’s hyperparameter settings, only modifying the learning rate and decoder depth. The hyperparameters were initially determined on ViT-Base and then directly applied to ViT-Small, ViT-Large, and ViT-Huge. Both CrossMAE and MAE are trained for 800 epochs. After pre-training, we evaluate the pre-trained models by fine-tuning them for image classification and instance segmentation. We provide implementation details and more experiments on different datasets and downstream tasks (iNaturalist (Van Horn et al., 2018), Places365 (Zhou et al., 2017), and ADE20K (Zhou et al., 2019)) in Section B.5.
@@ 4.1 ImageNet Classification
Results. As shown in Table 1, CrossMAE outperforms vanilla MAE using the same ViT-B encoder in terms of fine-tuning accuracy. This shows that replacing the self-attention with cross-attention does not degrade the downstream classification performance of the pre-trained model. Moreover, CrossMAE outperforms other self-supervised and masked image modeling baselines, e.g., DINO (Caron et al., 2021), MoCo v3 (Chen et al., 2021b), BEiT (Bao et al., 2021), and MultiMAE (Bachmann et al., 2022).
[…]
MAE 83.0
CrossMAE 83.3
CrossMAE + Self-Attn 83.3
(a) Attention type in decoder blocks. Adding back self-attention between mask tokens does not improve performance.
[…]
$65\%$ 83.5
$75\%$ 83.3
$85\%$ 83.3
(b) Mask ratio. CrossMAE has consistent performance across high mask ratios.
[…]
$15\%$ 83.1
$25\%$ 83.2
$75\%$ 83.3
(c) Prediction ratio. CrossMAE performs well even when only a fraction of mask tokens are reconstructed.
[…]
1 82.9
3 83.3
6 83.5
12 83.3
(d) Inter-block attention. A combination of six select encoder feature maps is best.
[…]
1 83.0
4 83.1
8 83.1
12 83.3
(e) Decoder depth. CrossMAE performance scales with decoder depth.
[…]
224 83.2
448 84.6
(f) Input resolution. CrossMAE scales to longer input sequences.
Table 3: Ablations on CrossMAE. We report fine-tuning performance on ImageNet-1K classification with 400 epochs (i.e., half of the full experiments) with ViT-B/16. MAE performance is reproduced using the official MAE code. Underline indicates the default setting for CrossMAE. Bold indicates the best hyperparameter among the tested ones. $1$ feature map fused (row 1, Table 3(d)) indicates using only the feature from the last encoder block. We use $25\%$ prediction ratio for both settings in Table 3(f) to accelerate training.
@@ 4.2 Object Detection and Instance Segmentation
Setup. We additionally evaluate models pretrained with CrossMAE for object detection and instance segmentation, which require deeper spatial understanding than ImageNet classification. Specifically, we follow ViTDet (Li et al., 2022b), a method that leverages a Vision Transformer backbone for object detection and instance segmentation. We report box AP for object detection and mask AP for instance segmentation on the COCO dataset (Lin et al., 2014), following MAE (He et al., 2022). We compare against supervised pre-training, MoCo-v3 (Chen et al., 2021b), BEiT (Bao et al., 2022), and MAE (He et al., 2022).
Results. As listed in Table 2, CrossMAE, with the default $75\%$ prediction ratio, performs better compared to these baselines, including vanilla MAE. This suggests that similar to MAE, CrossMAE performance on ImageNet positively correlates with instance segmentation. Additionally, CrossMAE’s downstream performance scales similarly to MAE as the model capacity increases from ViT-B to ViT-L. This observation also supports our hypothesis that partial reconstruction is suprisingly sufficient for learning dense visual representation.
@@ 4.3 Ablations
Cross-Attention vs Self-Attention. As shown in Table 3a, CrossMAE, with its cross-attention-only decoder, outperforms vanilla MAE in downstream tasks as noted in Section 4.1. Additionally, combining cross-attention with self-attention does not enhance fine-tuning performance, indicating that cross-attention alone is adequate for effective representation learning.
Mask Ratio and Prediction Ratio. In our experiments with different mask and prediction ratios (i.e., the ratio of mask tokens to all tokens and the ratio of reconstructed tokens to all tokens, respectively) (see Table 3b and Table 3c), we found that our method’s performance is not significantly affected by variations in the number of masked tokens. Notably, CrossMAE effectively learns representations by reconstructing as few as 15% of tokens, compared to the 100% required by vanilla MAE, with minimal impact on downstream fine-tuning performance, which shows that partial reconstruction is sufficient for effective representation learning.
Inter-block Attention. Our ablation study, detailed in Table 3d, explored the impact of varying the number of encoder feature maps in our inter-block attention mechanism. We found that using only the last feature map slightly lowers performance compared to using all 12. However, even a partial selection of feature maps improves CrossMAE’s performance, with the best results obtained using 6 feature maps. This indicates that CrossMAE does not require all features for optimal performance.
Decoder Depth. Table 3e shows that a 12-block decoder slightly improves performance compared to shallower ones. Remarkably, CrossMAE achieves similar results to MAE with just one decoder block, demonstrating its efficiency. Our experiments in Figure 7 that models with lower prediction ratios benefit more from deeper decoders.
@@
Input Resolution. We extend CrossMAE to longer token lengths by increasing the image resolution with constant patch size. Escalating the resolution from 224 to 448 increases the token length from 197 to 785, challenging the scalability of current approaches. Thus, we opt for a CrossMAE variant with a 25% prediction ratio. In Table 3f, we observe that the classification accuracy positively correlates with the input resolution, indicating that CrossMAE can scale to long input sequences.
4.4 Improving Training Throughput and Memory Utilization
In this section, we show a tangible improvement that CrossMAE provides: by implementing partial reconstruction and confining attention to interactions between masked and visible tokens, CrossMAE significantly improves pre-training efficiency compared to vanilla MAE.
Method (prediction ratio) MAE (0.75) CrossMAE (0.75) CrossMAE (0.25)
Accuracy (%) 83.0 83.3 83.2
Runtime (mins per epoch) 9.35 8.41 6.32
Memory (MB per GPU) 68386 57987 36805
Table 4: Comparison of MAE and CrossMAE with different prediction ratios. Each model is trained for 400 epochs.
Method Accuracy (%) Runtime (mins per epoch)
MAE 83.0 9.35
+Cross Attention (Section 3.2) 82.9 7.38
+Partial Reconstruction (Section 3.3) 83.2 6.32
+Inter-block Attention (Section 3.4) 83.3 8.41
Table 5: A structured ablation study of each contributing component of CrossMAE on a ViT-B architecture. Each model is trained for 400 epochs using 2 NVIDIA A100 80GB GPUs.
In Table 4, we present a comparison of ImageNet classification accuracy, runtime, and memory usage. In Table 5, we provide a more structured ablation of the different components of CrossMAE and their affect on runtime. The setup mirrors that of Table 3, with runtime measured on 2xA100 80GB GPUs, utilizing Flash-Attention 2 (Dao et al., 2023) across all models and gradient accumulation set to 2 for 400 epochs.
It is important to note that partial reconstruction relies on our proposed use of cross-attention decoder blocks, which is not applicable to vanilla MAE for efficiency improvements. In vanilla MAE, each masked token attends to other masked tokens within the decoder blocks. As a result, removing any masked tokens affects the decoding of the remaining ones. In contrast, CrossMAE reconstructs each masked token solely based on the visible tokens, without relying on interactions with other masked tokens. This allows the loss applied to the subset to serve as an unbiased estimate of the loss applied to all reconstructed tokens. This method reduces the computational load and runtime without sacrificing significant performance. Further discussion can be found in Appendix C.
Since vanilla MAE uses a self-attention decoder, the computational complexity of the decoder is quadratic with respect to the total number of tokens. Using cross-attention in CrossMAE reduces this complexity to be linear with respect to the number of visible tokens and the number of masked tokens. By further varying the number of masked tokens through partial reconstruction, we can reduce the complexity by decoding only a subset of the masked tokens.
Formally, let the total number of visible tokens be $N$, and the latent dimension be $d$, with a mask ratio $p$ and prediction ratio $\gamma$. The attention complexity in the MAE decoder is on the order of $N^{2}d$. In CrossMAE, the visible tokens (of length $(1-p)N$) serve as the queries, and the masked tokens (of length $\gamma N$) serve as the keys and values. This results in the attention complexity in CrossMAE decoder being $(1-p)\gamma N^{2}d$. For instance, in Table 1, the CrossMAE variant that decodes all masked tokens, or CrossMAE (0.75), uses $p=0.75$ and $\gamma=0.75$ (i.e., full prediction with $p=\gamma$), which gives $\frac{3}{16}N^{2}d$. On the other hand, CrossMAE (0.25) uses $p=0.75$ and $\gamma=0.25$, which gives $\frac{1}{16}N^{2}d$, further reducing the complexity. More discussion and comparisons can be found in Appendix C.
Figure 5: We visualize the output of each decoder block. (a-b) Different decoder blocks play different roles in the reconstruction, with most details emerging at later decoder blocks, which confirms the motivation for inter-block attention. (c) Visualizations of inter-block attention shows that different decoder blocks indeed attend to feature from different encoder blocks, with later blocks focusing on earlier encoder features to achieve reconstruction. The reconstructions are unnormalized w.r.t ground truth mean and std for each patch.